A physician advises a 220 lb man who is 6 feet tall to exercise at 10 METs for 10 minutes. If he follows the advice, how many kilocalories does he burn in 10 minutes?
Answer #1

ANSWER.

* Calculation of kilocalories (kcal) burn in 10 minute.

       Formula

         Calorie/hour = MET * weight in kg

Here

         MET = 10 minute.

         Weight = 220 lb

Convert lb to kg for calculation

             1 lb = 0.454 kg

        So

          Weight of the patient = 220 * 0.454

                                            = 99.8 kg.

   As per above formula

               Calorie/hour = 10 * 99.8

                                    = 998 calorie/hr.

           Calorie burn in 10 minute

                                 998 calorie = 60 min

                                      X calorie = 10 min

                            X * 60 = 998 * 10

                                    X = 9980 / 60

                                         = 166.3 calorie/10 minute

           Kcal burn in 10 minutes

                               1 calorie = 0.001 kcal

                   So

                        Kilocalorie burn in 10 min = 166.3 * 0.001

                                                                   = 0.17 kcal /10 minute.